OpenStudy (anonymous):

I am doing solving Quadratic equations by factoring and Applications of quadratic equations. I have a test TODAY! I dont understand it at all I am doing solving Quadratic equations by factoring and Applications of quadratic equations. I have a test TODAY! I dont understand it at all @Mathematics

OpenStudy (anonymous):

What is a quadratic equation?

OpenStudy (anonymous):

hold on I will give you one to look at

OpenStudy (anonymous):

y^2 + 3y +2 = 0

OpenStudy (anonymous):

\[Ax^2 + Bx + C = 0\]

OpenStudy (anonymous):

It says solve the equation and check your solutions I just dont understand this

OpenStudy (anonymous):

There are so many applications of quadratic equations, but solving them by simply factoring might require a bit of.... 'intuition'. can you explain how did I rewrite \[y^2 + 3y + 2 = 0\]as\[(y+2)(y+1)=0\]

OpenStudy (anonymous):

I am just so confused on how to solve this and my teacher goes so fast that my brain can't absorb it

OpenStudy (anonymous):

but she brakes it all down and says y +2= something and y - 1= 0

OpenStudy (anonymous):

do you know how to factor quadratic expressions?

OpenStudy (anonymous):

not really lol. Im sorry. I know I am confusing but its bc I am so confused

OpenStudy (anonymous):

I just need someone to break down the steps to solve the problems. Like dummify them I guess

OpenStudy (anonymous):

first step to solving a quadratic equation 'by factoring' is to factor the quadratic expression \[Ax^2 + Bx + C\] Look for two numbers that multiply together to equal C, and add up to equal B. for example, for \[y^2 + 3y + 2\] notice that the numbers 2 and 1 multiply to 2 and add up to 3

OpenStudy (anonymous):

once you've found those 2 numbers, you can factor the quadratic expression \[y^2+3y+2\] into \[(y+2)(y+1)\]

OpenStudy (anonymous):

Basically, factorise the equation you have, then make it equal to 0: (y+2)(y+1)=0 Then take each equation and work out x, so: (y+2)=0, so y must equal -2 and (y+1)=0, so y must equal -1. Those are your two x-coordinates.

OpenStudy (anonymous):

Ok. I got that part down. so I guess I do know the factoring part lol

OpenStudy (anonymous):

ok so basically what you are saying is change the signs like (y + 2) to -2. and +1 to -1 and Im done?

OpenStudy (anonymous):

sorry had a phone call once you've factored the quadratic expression, it becomes very easy to solve the quadratic equation \[(y+2)(y+1)=0\] the equation is now telling you that either factor, y +2, or y +1, can equal to zero, so you can consider 2 possible solutions \[y+2 = 0\] \[y + 1 = 0\] solve those 2 equations to get \[y = -2\ \ \ or\ \ \ y = -1 \]

OpenStudy (anonymous):

Yes, in a sense. You've just got to make y+2=0 and y+1=0

OpenStudy (anonymous):

ok heres another one. (6k + 5) (k+4) = 0 some how she made it into a fraction 6k/6 = 5/6 and the answer ended up being k+4=0 k =-4 How did that happen? lol

OpenStudy (anonymous):

it's the same thing\[(6k+5)(k+4)=0\] you can break the above equation down to two equations:\[6k+5 = 0\]\[k + 4 = 0\] solving both of those equations gives us (I'll start with the first one)\[6k = -5\]\[k = -\frac{5}{6}\] and solving the second equation gives us \[k = -4\] so our solutions are y = -5/6 or y = -4

OpenStudy (anonymous):

Also I have a big problem with knowing if its supposed to be a positive or negative, Like when your doing the foil method on a problem with a positive and 2 negs.

OpenStudy (anonymous):

what's the foil method? :(

OpenStudy (anonymous):

First Outter Inner Last

OpenStudy (anonymous):

I've never heard of that method. I might have used it before though :-P

OpenStudy (anonymous):

but how did you get -5 from 6k?

OpenStudy (anonymous):

solve this equation \[6k + 5 = 0\]

OpenStudy (anonymous):

well basically you muliply the first of the equation times the first of the second... (y + 2) (4y- 1) you would multiply y * 4

OpenStudy (anonymous):

would it be -1?

OpenStudy (anonymous):

k= -1 i mean

OpenStudy (anonymous):

how did you get k = -1?

OpenStudy (anonymous):

6k - 5? was that wrong?

OpenStudy (anonymous):

For that one you just wrote, @bhunsuck, the foiled answer is: 4y^2+-1y+8y-2. This'll give you: 4y^2+9y-2

OpenStudy (anonymous):

oh! you mean expanding the expression :-P

OpenStudy (anonymous):

but is the problem done?

OpenStudy (anonymous):

I'm sure it is :-D

OpenStudy (anonymous):

so the answer will be 4y^2+9y-2=0

OpenStudy (anonymous):

For (y + 2) (4y- 1), yes it is.

OpenStudy (anonymous):

huh.. cool... do you mind working a few more problems? I do better with repetition

OpenStudy (anonymous):

Yeah, sure. Fire away. :)

OpenStudy (anonymous):

ty :D

OpenStudy (anonymous):

6\[x^2\]= 4+\[5x\]

OpenStudy (anonymous):

ooops sorry, im not very good with this thing yet lol

OpenStudy (anonymous):

you will get better :-D

OpenStudy (anonymous):

Is that \[6x ^{2}= 4+ 5x\]?

OpenStudy (anonymous):

i know you have to change te signs and go from squares, variable then numbers such as 6x^2 - 5x-4=0

OpenStudy (anonymous):

yes it is damned

OpenStudy (anonymous):

hello lol

OpenStudy (anonymous):

\[6x ^{2}-5x-4=0\] Then, factorise: (\[\left( -2x-1 \right)\left( -3x+4 \right)\] Then, make both equations equal to zero and figure them out, so: (-2x-1)=0 and (-3x+4)=0 Finally, find out x, which seems pretty difficult. I think it's: x=1/-2 and x=-4/-3

OpenStudy (anonymous):

...

OpenStudy (anonymous):

ok so how did u get the -2x -1 and -3 +4?

OpenStudy (anonymous):

Sorry I took so long. Was trying to figure out the last bit. Just factorise. It's basically trial and error and you have to try different numbers until you figure it out. For this one, I knew the two x's would either be 6x and x or 3x and 2x because they multiply to make 6x^2. Then I had to find two other numbers that multiplied to make the last number but added up to make the middle number, but taking the x number's into account. This may be a slightly terrible way of describing it to you XD

OpenStudy (anonymous):

yeah kinda is hard to understand lol

OpenStudy (anonymous):

ur answer x = -1/2 and x= 4/3 is right btw

OpenStudy (anonymous):

Okay, what don't you get, specifically? Thank god for that XD

OpenStudy (anonymous):

Oh jesus... I know how to factor.. I know I know how to factor... But when I look at these problems it all goes out of my head! lol hold on a sec. I am trying to soak this in and figure it out.

OpenStudy (anonymous):

I know how it feels, doing revision just before a test. Didn't go to badly, luckily.

OpenStudy (anonymous):

OMG, ok lol... My husband is soooooo right.. (shhh I didn't say that) I over think things and then I confuse myself! Ok I got the first part. I feel so silly lol

OpenStudy (anonymous):

I was lost when I first did this stuff too :-P

OpenStudy (anonymous):

hi agdg, wb :)

OpenStudy (anonymous):

Haha! You have a sensible husband then. Okay, what other bit don't you get?

OpenStudy (anonymous):

ok so let me get this straight... U got -2x -1 and -3 +4 bc 2x 3= the 6 and added together makes the 5. and 1 x 4 = 4 and added together makes the 5 right?

OpenStudy (anonymous):

How do you know to turn it into a fraction? when theres more than one number in the answer?

OpenStudy (anonymous):

XD Kinda. I got the -2x and -3x because yes they do make 6 but then I had to pick numbers that added up to 5 but multiplied to make 4, taking into account that those two numbers had to be multiplied by the 3x and 2x. Well, when I got the two brackets, all I did was rearrange them to make it equal to x, so: (-2x-1)=0 -2x=1 x=1/-2. You remove the brackets by the way because they're not really serving any purpose.

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

can u explain x^2=7x? The way I think is just take the x-7=0 x=7

OpenStudy (anonymous):

Divide everything by x and you have your answer for that one. So, yes, x=7

OpenStudy (anonymous):

so y in gods name dont they just say take out the x's and that will be your answer? lol

OpenStudy (anonymous):

be careful about dividing by x: what if x = 0?

OpenStudy (anonymous):

well how would i know? My main goal at this point is to have to learn as LITTLE steps to solve as possible bc my brain can't hold all the different crap

OpenStudy (anonymous):

instead of dividing, try taking the square root of both sides :-D

OpenStudy (anonymous):

i dont know how to find the square roots. thats one of mt problems

OpenStudy (anonymous):

you're right; the square root won't work for that particular equation :( go ahead and divide by x, but keep in mind that one of the possible solutions is x = 0

OpenStudy (anonymous):

But you can't really square root 0, can you?

OpenStudy (anonymous):

i can find them in the smaller numbers but not the bigger ones.

OpenStudy (anonymous):

the square root of zero is zero. still doesn't work with the equation x^2 = 7x though :( I would divide by x x = 0 or x = 7

OpenStudy (anonymous):

ok here one where y = 0 but I dont understand how she got that answer

OpenStudy (anonymous):

what's the equation?

OpenStudy (anonymous):

10y^2 = 5y

OpenStudy (anonymous):

same reason: one possible solution to 10y^2 = 5y is y = 0

OpenStudy (anonymous):

well we are supposed to list both solutions. y=0 y= 1/2

OpenStudy (anonymous):

I have like 30 problems to do before class in like 2 hours. So i guess I better go. THANK YOU SO MUCH for your help. Do you come on here alot? I know I will need more help lol